OpenStudy (anonymous):

Purvi and her friends tipped their waiter a total of ten dollars for dinners that cost nine dollars, eight dollars, eleven dollars and twelve dollars each. Use a percent proportion to solve for the percent tip they left the waiter. Show your work.

OpenStudy (anonymous):

x/100 = $12/($9+$8+$11+$12) 1200=40*x x=30% olive oil=75% or 0.75 3 = 0.75*x x=4 (total solution volume is 4 cups) $80,000*0.15=$12,000 $12,000+$40,000=$52,000 $58,000 is more than $52,000 so Isabelle made more money than Randall. 143-130=13 13 = 0.01*x*130 13 = 1.3*x 10 = x (10% increase)

OpenStudy (anonymous):

1. The total cost all dinners was $9 + $8 + $11 + $12 = $40. The amount of the tip is $12. Thus, they tipped a total of $12/$40 = 30%. That's a big tip! 2. 25% is 1/4, and 75% is 3/4. Therefore, to make 4 cups of furniture polish, you'd mix 1 cup of vinegar with 3 cups of olive oil. 3. Randall's income was $40,000 + 0.15 × $80,000 = $40,000 + $12,000 = $52,000. Isabelle's income of $58,000 was more. 4. The percent increase is the difference divided by the original price. Thus, the percent increase is ($143 - $130)/$130 = $13/$130 = 0.10 = 10%

OpenStudy (anonymous):

x/100 = $12/($9+$8+$11+$12) 1200=40*x x=30%